OpenStudy (lgbasallote):

Each molecule of ampicillin contains 16 atoms of carbon (plus other atoms). The mass percent of carbon in ampicillin is 55.00%. What is the molar mass of ampicillin? A) 698.8 g/mol B) 105.7 g/mol C) 192.2 g/mol D) 290.9 g/mol E) 349.4 g/mol do i divide the mass of carbon by the mass percent?

OpenStudy (xishem):

If you've got 16 atoms of carbon, you can convert that to some mass of carbon. If you know that that mass is 55.0% of the total compound, you can set up a ratio:\[\frac{55.0\%}{100.0\%}=\frac{x\ g}{Total\ g}\]That will give you the total mass of one molecule of the compound, and since the molar mass is the same thing as the mass per molecule, you've got your answer.

OpenStudy (lgbasallote):

so is it just dividing mass of carbon by the mass percent?

OpenStudy (xishem):

The way you're wording it seems strange. The percent composition of a substance is defined as the % (by mass) composition of each element of the compound. If I have 10g of carbon in a 100g substance, the percent composition of that compound in respect to carbon is 10%.

OpenStudy (lgbasallote):

you do know you still havent really answered whether it's mass of carbon divided by mass % or not...

OpenStudy (lgbasallote):

what does 16 atoms of C mean @Xishem

OpenStudy (lgbasallote):

hmm is it \[\huge \frac{16\times 12.01}{55} \times 100\]

OpenStudy (lgbasallote):

because it seems i got the answer using that formula

OpenStudy (xishem):

16 atoms of C means 16 atoms of carbon. "it's mass of carbon divided by mass % or not..." What is "it" in this case? I've told you what the definition of percent composition is.

OpenStudy (lgbasallote):

just want to know if it's a fluke or is that really how to solve it

OpenStudy (xishem):

Well, let's break what you did down.

OpenStudy (lgbasallote):

...i just want to know if the formula is a fluke or not...

OpenStudy (lgbasallote):

@Xishem is the formula wrong or not??

OpenStudy (xishem):

You took 16 atoms of carbon and converted it to a mass of carbon...\[16\ atoms\ C \times \frac{12.01g\ C}{1\ atom\ C}=192.16g\ C\]Then you divide by the percent composition of the compound in respect to carbon... So if we look back to that relation I described earlier:\[\frac{55.0\%}{100.0\%}=\frac{x\ g}{Total\ g} \rightarrow Total\ g=\frac{x\ g}{50\%}\times 100\%\]So, yes. That formula is correct. It can be related back to the relation I gave to you earlier.

OpenStudy (lgbasallote):

ahh finally. a direct answer.

OpenStudy (xishem):

It's more important that you understand where the relationship comes from than that I give you a direct answer.

OpenStudy (lgbasallote):

well i was asking for a yes or no and you kept on being indirect....

OpenStudy (lgbasallote):

i had so many ideas in mind i just dont know which one...

OpenStudy (xishem):

The thing is that taking the mass of carbon divided by the "percent composition" of carbon in the molecule multiplied by 100% doesn't make sense intuitively. You can try to memorize that that's how you convert from a mass of carbon and percent composition to the molar mass of the compound, but that's a really ineffective way to learn. A lot of quantitative sciences revolve around knowing how to work through units without knowing much about the equations involved. Tons and tons of equations and relationships can be derived by simply working through the units.